Assume it is May 2018. A company expects to receive a cash payment of £10

million in 7 months’ time from a customer. The company will not have immediate use

of the cash until March next year. During this idling period, the financial manager will

deposit this amount of cash for three months to earn interest. The price for the

December 3-month interest rate futures in May 2018 is 94. The financial manager is

concerned the interest rate will fall when she receives the cash.

a) Design a strategy using short-term interest rate futures to hedge against the

company’s exposure to the risk of changing the interest rate. Suppose the size

of interest-rate futures is £500,000.

b) Suppose the interest rate falls to 3.1% in December 2018. By how much is the

amount which the company receives less than the interest from the deposit

(compared to the level of interest in May 2018). In other words, what is the

loss incurred by the company regarding interest rate income in the cash

market?

c) Calculate the gain from the futures market in December 2018 at expiry.
Answer #1

a) the firm is investor and it will like to off set interest rate loss by profit on futures contract. Accordingly if Interest rate falls it will gains so it will take long position in future contract. That means he should buy interest rate futures.

No. Of contracts = Amount of investment/size contract

= 10000000/500000

= 20 contracts

(b)

Actual Interest rate loss = 6%(100-94) -3.10% = 2.9%

Amount of loss = 10000000*2.9%*3/12 = 72500

loss in cash market = 72500

(c)

We have taken long position in Interest rate futures at 94.

What will happen in future

Buy to open = 94

Sell to close = 96.9(100-3.10)

Gain from future = 96.9-94 = 2.9%

Amount = 10000000*2.9%*3/12 = 72500

so gain in future Market will be = 72500